My sister is 14 years older than I am. The sum of our ages is 28. Find our ages

Answer by ikleyn(52794)

.

From  28  years,  subtract the difference of ages of  14  years.

The difference,  28-14 = 14,  is your doubled age.

Hence,  your age is  14/2 = 7 years.

Then your sister is  7 + 14 = 21 years  old.

-------------

Solved mentally,  without using equations.